To cyclist travelling due north along a straight road at 10km/h,the wind appears to come from the east.When he increases his speed to 15km/h,the wind appears to come from the direction N67°E.Find the speed and the direction of the wind over the ground.


1
Expert's answer
2021-05-07T09:19:46-0400

The initial velocity of cyclist is


"\vec V_{C1}=10\vec j"

The wind appears to come from the east.

Suppose the relative velocity of the wind is


"\vec V_{WC_1}=-a\vec i+0\vec j, a>0"

The velocity of the wind is


"\vec V_W=\vec V_{C1}+\vec V_{WC_1}=-a\vec i+10\vec j"


 Cyclist increases his speed to 15km/h


"\vec V_{C2}=15\vec j"

The wind appears to come from the direction N67°E.

Suppose the relative velocity of the wind is


"\vec V_{WC2}=-c\sin(67\degree)\vec i-c\cos(67\degree)\vec j, c>0, d>0"


The velocity of the wind is


"\vec V_W=\vec V_{C2}+\vec V_{WC_2}"


"=-c\sin(67\degree)\vec i+(15-c\cos(67\degree))\vec j"

Then


"\begin{matrix}
 a =c\sin(67\degree) \\
 10=15-c\cos(67\degree)
\end{matrix}"

"c=\dfrac{5}{\cos(67\degree)}"

"a=5\tan(67\degree)"

"\vec V_W=-5\tan(67\degree)\vec i+10\vec j"

"|\vec V_W|=\sqrt{(-5\tan(67\degree))^2+(10)^2}"

"=5\sqrt{4+\tan^2(67\degree)}"

"\tan(\theta)=\dfrac{10}{-5\tan(67\degree)}"

"\theta=180\degree-\tan^{-1}(\dfrac{2}{\tan(67\degree)})"


"|\vec V_W|\approx15.45\ km/h"

"\theta\approx139.67\degree"